# On the Wisdom of Crowds

### v2

@@
Abstract: We study the properties of macroeconomic survey forecast response averages as the number of survey respondents grows. Such averages are “portfolios” of forecasts. We characterize the speed and pattern of the gains from diversification as a function of portfolio size (the number of survey respondents) in both (1) the key real-world data-based environment of the U.S. Survey of Professional Forecasters, and (2) the theoretical model-based environment of equicorrelated forecast errors. We proceed by proposing and comparing various direct and model-based “crowd size signature plots”, which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We then estimate the equicorrelation model for growth and inflation forecast errors by choosing model parameters to minimize the divergence between direct and model-based signature plots. The results indicate near-perfect equicorrelation model fit for both growth and inflation, which we explicate by showing analytically that, under very weak conditions, the direct and fitted equicorrelation model-based signature plots are identical at a particular model parameter configuration. That parameter configuration immediately suggests an analytic closed-form estimator for the direct signature plot, so that equicorrelation ultimately emerges as a device for convenient calculation of direct signature plots, rather than a separate “model” producing separate signature plots. In any event we find that the gains from survey diversification are greater for inflation forecasts than for growth forecasts, and that they are largely exhausted with inclusion of 5-10 representative forecasters.
@@ 1 Introduction and Basic Framework
The wisdom of crowds, or lack thereof, is traditionally and presently a central issue in psychology, history, and political science; see for example Surowiecki (2005) regarding wisdom, and Aliber et al. (2023) regarding lack thereof. Perhaps most prominently, however, the wisdom of crowds is also—and again, traditionally and presently—a central issue in economics and finance, where heterogeneous information and expectations formation take center stage. [^1] Interestingly, moreover, it also features prominently in new disciplines like machine learning and artificial intelligence, via forecast combination methods like ensemble averaging (e.g., Diebold et al., 2023).
@@
In particular, we study SPF crowd behavior as crowd size grows, asking precisely the same sorts of questions of SPF “forecast portfolios” that one asks of financial asset portfolios: [^4] Classic early work on which we build includes, for example, Makridakis and Winkler (1983) and Batchelor and Dua (1995). We progress much farther, however, particularly as regards analytic characterization. How quickly, and with what patterns, do diversification benefits become operative, and eventually dissipate, as portfolio size (the number of forecasters) grows, and why? Do the results differ across the key SPF variables (growth and inflation), and if so, why? What are the implications for survey size and design?
@@
We answer the above questions using what we call “crowd size signature plots”, which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We examine both “direct signature plots” (estimated directly from the SPF data) and “model-based signature plots” (estimated by fitting a simple model of forecast-error equicorrelation), and we characterize their surprisingly-intimate relationship, uncovering en route a remarkably simple analytic closed-form signature plot estimator. We focus throughout on estimating and interpreting signature plots for both growth and inflation, providing individual and comparative assessments of their paths and patterns.
@@ 2 Direct SPF Crowd Size Signature Plots
      ($\widehat{MSE}^{avg}_{R,NT}(k)$ crowd size signature plots”, where “RR” denotes “relative”, or “ratio”);
@@
      ($\widehat{DMSE}^{avg}_{NT}(k)$ crowd size signature plots”, where “DD” denotes “delta”, or “change”),
@@
$\widehat{MSE}^{avg}_{NT}(k)$ and the related objects above are are readily calculated in principle, but complications arise in practice. In particular, calculation is infeasible unless $N$ is very small, due to the huge number of different $k$-average forecasts. For example, for $N=40$ and $k=20$ we obtain ${N\choose k}=O(10^{11})$. Hence we proceed by approximating $\widehat{MSE}^{avg}_{NT}(k)$ as follows:
@@
The growth vs inflation $\widehat{MSE}^{avg}_{R,NT}(k)$ plots asymptote to very different levels as $k$ increases – approximately 80% for growth and 60% for inflation – which highlights a central result: The benefits of SPF “portfolio diversification” appear to be substantially greater for inflation than for growth.
@@ 3.1 Population Results Under Equicorrelation
(Notice that $\sigma$ cancels in the $R^{avg}_{N\infty}(k;\sigma)$ calculation, so we simply write $R^{avg}_{N\infty}(k)$.)
@@
Importantly, the optimality of simple averaging under zero correlation is preserved under equicorrelation. That is, equicorrelation is sufficient for the optimality of simple averaging – an immediate implication of the results of Elliott (2011), who shows that a necessary and sufficient condition for optimality of simple averaging is that row sums of $\Sigma$ be equal. Equicorrelation is one such case, although there are of course others, obtained by manipulating correlations in their relation to variances to keep row sums equal, but none are nearly so compelling and readily interpretable as equicorrelation.
@@
Overall, the value of averaging across forecasters is determined by the interaction of $k$ and $\rho$. Other things equal, it is highest when $k$ is small (small pool) or when $\rho$ is low (weakly-correlated forecast errors). In particular, for realistic values of $\rho$, around 0.5, say, most gains from increasing $k$ are obtained by $k=5$.
@@
The fact that, for realistic values of $\rho$, most $MSE^{avg}_{R,N\infty}(k;\rho)$ gains from increasing $k$ are obtained by $k=5$ does not necessarily indicate that typical surveys use too many forecasters. $MSE^{avg}_{R,N\infty}(k;\rho)$ is an average across all $k$-forecast combinations, and the best and worst $k$-average combinations, for example, will have very different MSEs. Figure 6 speaks to this; it shows $MSE^{min}_{R,N\infty}(k;\rho)$, $MSE^{avg}_{R,N\infty}(k;\rho)$, and $MSE^{max}_{R,N\infty}(k;\rho)$) under equicorrelation with $\rho=0.5$, for k=1,,20k=1,...,20.101010We use $N=40$ as an approximation to the average number of forecasters participating in the SPF in any given quarter, and we use $T=160$ to mimic the total sample size when working with 40 years of quarterly data, as in the SPF.
@@ 3.2 Estimating the Equicorrelation Model
Here we estimate the equicorrelation model by choosing its parameters $\rho$ and $\sigma$ to make the equicorrelation $MSE^{avg}_{N\infty}(k;\rho,\sigma)$ as close as possible to the SPF $\widehat{MSE}^{avg}_{NT}(k)$. This estimation strategy is closely related to, but different from, GMM estimation. Rather than matching model and data moments, it matches more interesting and interpretable functions of those moments, namely model-based and direct crowd size signature plots – as per the “indirect inference” of Smith Jr (1993) and Gourieroux et al. (1993). Henceforth we refer to it simply as the “matching estimator”.
@@ 3.2.1 Calculating the Solution
Calculating the minimum in equation (21) is very simple, because the bivariate minimization can be reduced to a univariate mimimization in $\rho\in\left]\frac{-1}{N-1},1\right[$. To see this, recall that $MSE_{N\infty}^{avg}(k;\rho,\sigma)=\frac{\sigma^{2}}{k}[1+(k-1)\rho]$, so that the first-order condition for $\sigma^{2}$ is
@@ 3.2.2 Understanding the Near-Perfect Equicorrelation Fit
Combining Remarks 13 and 14, it emerges that, regardless of whether one chooses to approach the signature plot from the “direct side” or from the “equicorrelation side”, there is no need for simulation (as earlier from the direct side, when implementing the direct estimator via Algorithm 1) or numerical optimization (as earlier from the equicorrelation side, when implementing the matching estimator via solution of the first-order conditions (22)-(23)). Instead, the trivially-simple closed-form estimator (29)-(30) is always applicable.
@@
We earlier observed that the empirical $DMSE^{avg}_{R,N\infty}(k)$ signature plots for growth and inflation appeared identical, and the corollary shows that indeed they must be identical.
@@ 3.3 On the Assessment of Equicorrelation Model Fit
As we have seen, equicorrelation ultimately emerges as a device for convenient calculation and understanding of direct signature plots, rather than as a separate “model” producing separate signature plots. In particular, nothing we have done requires that the the equicorrelation be “true”. Nevertheless, it may be of interest in some contexts to assess whether forecast errors are truly equicorrelated, which can be done (under stronger assumptions than those invoked thus far) by maximum-likelihood estimation of a dynamic-factor model, followed by likelihood-ratio tests of the equicorrelation restrictions.
@@
The DFM (31) is already in state-space form, and one pass of the Kalman filter yields the innovations needed for exact Gaussian likelihood evaluation and optimization, and it also accounts for missing observations associated with survey entry and exit. One may also impose the weak or strong equicorrelation restrictions (33) or (34), respectively, and assess them using likelihood-ratio tests. Unsurprisingly, such formal tests (not reported) reject equicorrelation – a stylized “toy” DGP if ever there was one – for both growth and inflation, so we will not pursue formal testing.
@@ 4 Summary, Conclusions, and Directions for Future Research
We have studied the properties of macroeconomic survey forecast response averages as the number of survey respondents grows, characterizing the speed and pattern of the “gains from diversification” and their eventual decrease with “portfolio size” (the number of survey respondents) in both (1) the key real-world data-based environment of the U.S. Survey of Professional Forecasters (SPF), and (2) the theoretical model-based environment of equicorrelated forecast errors. We proceeded by proposing and comparing various direct and model-based “crowd size signature plots”, which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We then estimated the equicorrelation model for growth and inflation forecast errors by choosing model parameters to minimize the divergence between direct and model-based signature plots.

### v3

@@
Abstract: We study the properties of macroeconomic survey forecast response averages as the number of survey respondents grows. Such averages are “portfolios” of forecasts. We characterize the speed and pattern of the gains from diversification as a function of portfolio size (the number of survey respondents) in both (1) the key real-world data-based environment of the U.S. Survey of Professional Forecasters, and (2) the theoretical model-based environment of equicorrelated forecast errors. We proceed by proposing and comparing various direct and model-based “crowd size signature plots,” which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We then estimate the equicorrelation model for growth and inflation forecast errors by choosing model parameters to minimize the divergence between direct and model-based signature plots. The results indicate near-perfect equicorrelation model fit for both growth and inflation, which we explicate by showing analytically that, under very weak conditions, the direct and fitted equicorrelation model-based signature plots are identical at a particular model parameter configuration. That parameter configuration immediately suggests an analytic closed-form estimator for the direct signature plot, so that equicorrelation ultimately emerges as a device for convenient calculation of direct signature plots, rather than a separate “model” producing separate signature plots. Finally, we find that the gains from survey diversification are greater for inflation forecasts than for growth forecasts, and that they are largely exhausted with inclusion of 5-10 representative forecasters.
@@
Disclaimer:: The views expressed here are those of the authors and do not necessarily represent those of the Federal Reserve Bank of Philadelphia or the Federal Reserve System.
@@ 1 Introduction and Basic Framework
The wisdom of crowds, or lack thereof, is traditionally and presently a central issue in psychology, history, and political science; see for example Surowiecki (2005) regarding wisdom, and Aliber et al. (2023) regarding lack thereof. Perhaps most prominently, however, the wisdom of crowds is also—and again, traditionally and presently—a central issue in economics and finance, where heterogeneous information and expectations formation play a prominent role. [^1] Interestingly, moreover, it also features prominently in new disciplines like machine learning and artificial intelligence, via forecast combination methods like ensemble averaging (e.g., Diebold et al., 2023).
@@
In particular, we study SPF crowd behavior as crowd size grows, asking precisely the same sorts of questions of SPF “forecast portfolios” that one asks of financial asset portfolios: [^4] Classic early work on which we build includes, for example, Makridakis and Winkler (1983) and Batchelor and Dua (1995). We progress much further, however, particularly regarding analytic characterization. How quickly, and with what patterns, do diversification benefits become operative, and eventually dissipate, as portfolio size (the number of forecasters) grows, and why? Do the results differ across the key SPF variables (growth and inflation), and if so, why? What are the implications for survey size and design?
@@
We answer the above questions using what we call “crowd size signature plots,” which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We examine both “direct signature plots” (estimated directly from the SPF data) and “model-based signature plots” (estimated by fitting a simple model of forecast-error equicorrelation), and we characterize their surprisingly intimate relationship, uncovering en route a remarkably simple analytic closed-form signature plot estimator. We focus throughout on estimating and interpreting signature plots for both growth and inflation, providing individual and comparative assessments of their paths and patterns.
Related Literature.
This paper contributes to the forecast combination literature by examining the performance patterns of simple $k$-average forecasts as the number of forecasters, $k$, increases. We note that this paper is not focused on the problem of optimal forecast combination (see Bates and Granger (1969) for a pioneering contribution, and for example Timmermann (2006), Hsiao and Wan (2014) and Wang, Hyndman, Li, and Kang (2022) for recent developments). The motivation for focusing on simple $k$-averages is twofold. First, it is a well-established empirical fact that simple averages of forecasts produce combined forecasts with surprisingly good forecasting performance despite not being optimal in general (e.g., Clemen (1989), Stock and Watson (2004), and Genre, Kenny, Meyler, and Timmermann (2013)). [^5] This empirical fact has been called a “forecast combination puzzle”; work focused on studying the causes of this puzzle includes Smith and Wallis (2009), Elliott and Liao (2025), and Claeskens, Magnus, Vasnev, and Wang (2016). Second, simple averages can be optimal in specific contexts. For example, Engle and Kelly (2012) show that when forecast errors exhibit equicorrelation, equal weights are optimal; Elliott and Liao (2025) later provide precise necessary and sufficient conditions under which simple averages are optimal.
Our focus on performance patterns of simple k-averages as k increases builds on earlier empirical work. Makridakis and Winkler (1983), using the M-Competition time-series dataset and combinations of forecasts across multiple forecasting methods, document that simple-average combinations become more accurate, with sharply diminishing marginal gains, as the number of component forecasts increases. We find the same qualitative pattern in the SPF when averaging across professional forecasters. We extend these results by providing a theoretical foundation: we derive an exact expression for the direct crowd size signature plot that links its level and slope to two interpretable features of SPF forecast errors (average dispersion and average dependence). This delivers an equivalent equicorrelation representation and explains both the hyperbolic shape of the signature plots and the rapid exhaustion of diversification gains.
In another related paper, Chan and Pauwels (2018) propose a framework to study the properties of a combined forecast, and derive conditions under which the simple average is optimal. In their setting, they point out that the performance of a combined forecast improves as the number of forecasts or models increases, and that a simple average will asymptotically outperform any fixed individual forecast. Although this is consistent with our results, our emphasis is more general, as we characterize the entire trajectory of performance improvements, including the practically relevant setting of a small to moderate number of forecasters.
Our results on relating direct crowd size signature plots to an equicorrelation model-based signature plot are connected to the results in Clemen and Winkler (1985). In the case of normally distributed forecast errors, they show that the value of additional information for forecast combination diminishes when the additional information is correlated. This is consistent with our findings documenting diversification gains; however, our results go beyond this by not requiring the normality assumption.
Finally, we contribute to the SPF survey-forecast literature (e.g., Croushore and Stark, 2019; Clements et al., 2022) by providing evidence and theory directly relevant for survey design: how many forecasters are needed to capture most of the benefits of averaging, and why do those benefits differ across variables? Using SPF growth and inflation forecasts, we document a steeply diminishing-returns pattern in crowd-size performance, with most improvements realized by about 5-10 representative forecasters. Moreover, we show that diversification gains are considerably larger for inflation than for growth, and we provide an exact theoretical characterization linking the inflation-growth difference to the underlying variance-covariance structure of individual forecast errors.
@@ 2 Direct SPF Crowd Size Signature Plots
      ($\widehat{MSE}^{avg}_{R,NT}(k)$ crowd size signature plots,” where “RR” denotes “relative,” or “ratio”);
@@
      ($\widehat{DMSE}^{avg}_{NT}(k)$ crowd size signature plots,” where “DD” denotes “delta,” or “change”),
@@
$\widehat{MSE}^{avg}_{NT}(k)$ and the related objects above are readily calculated in principle, but complications arise in practice. In particular, calculation is infeasible unless $N$ is very small, due to the huge number of different $k$-average forecasts. For example, for $N=40$ and $k=20$ we obtain ${N\choose k}=O(10^{11})$. Hence we proceed by approximating $\widehat{MSE}^{avg}_{NT}(k)$ as follows:
@@
The growth vs. inflation $\widehat{MSE}^{avg}_{R,NT}(k)$ plots asymptote to very different levels as $k$ increases – approximately 80% for growth and 60% for inflation – which highlights a central result: The benefits of SPF “portfolio diversification” appear to be substantially greater for inflation than for growth.
@@ 3.1 Population Results Under Equicorrelation
Notice that $\sigma$ cancels in the $R^{avg}_{N\infty}(k;\sigma)$ calculation, so we simply write $R^{avg}_{N\infty}(k)$.
@@
Importantly, the optimality of simple averaging under zero correlation is preserved under equicorrelation. That is, equicorrelation is sufficient for the optimality of simple averaging – an immediate implication of the results of Elliott and Liao (2025), who shows that a necessary and sufficient condition for optimality of simple averaging is that row sums of $\Sigma$ be equal. Equicorrelation is one such case, although there are of course others, obtained by manipulating correlations in their relation to variances to keep row sums equal, but none are nearly so compelling and readily interpretable as equicorrelation.
@@
Overall, the value of averaging across forecasters is determined by the interaction of $k$ and $\rho$. Other things equal, it is highest when $k$ is small (small pool) or when $\rho$ is low (weakly correlated forecast errors). In particular, for realistic values of $\rho$, around 0.5, say, most gains from increasing $k$ are obtained by $k=5$.
@@
The fact that, for realistic values of $\rho$, most $MSE^{avg}_{R,N\infty}(k;\rho)$ gains from increasing $k$ are obtained by $k=5$ does not necessarily indicate that typical surveys use too many forecasters. $MSE^{avg}_{R,N\infty}(k;\rho)$ is an average across all $k$-forecast combinations, and the best and worst $k$-average combinations, for example, will have very different MSEs. Figure 6 speaks to this; it shows $MSE^{min}_{R,N\infty}(k;\rho)$, $MSE^{avg}_{R,N\infty}(k;\rho)$, and $MSE^{max}_{R,N\infty}(k;\rho)$) under equicorrelation with $\rho=0.5$, for k=1,,20k=1,...,20.111111We use $N=40$ as an approximation to the average number of forecasters participating in the SPF in any given quarter, and we use $T=160$ to mimic the total sample size when working with 40 years of quarterly data, as in the SPF.
@@ 3.2 Estimating the Equicorrelation Model
Here we estimate the equicorrelation model by choosing its parameters $\rho$ and $\sigma$ to make the equicorrelation $MSE^{avg}_{N\infty}(k;\rho,\sigma)$ as close as possible to the SPF $\widehat{MSE}^{avg}_{NT}(k)$. This estimation strategy is closely related to, but different from, GMM estimation. Rather than matching model and data moments, it matches more interesting and interpretable functions of those moments, namely model-based and direct crowd size signature plots – as per the “indirect inference” of Smith Jr (1993) and Gourieroux et al. (1993). Henceforth we refer to it simply as the “matching estimator.”
@@ 3.2.1 Calculating the Solution
Calculating the minimum in equation (21) is very simple, because the bivariate minimization can be reduced to a univariate minimization in $\rho\in\left]\frac{-1}{N-1},1\right[$. To see this, recall that $MSE_{N\infty}^{avg}(k;\rho,\sigma)=\frac{\sigma^{2}}{k}[1+(k-1)\rho]$, so that the first-order condition for $\sigma^{2}$ is
@@ 3.2.2 Understanding the Near-Perfect Equicorrelation Fit
Combining Remarks 13 and 14, it emerges that, regardless of whether one chooses to approach the signature plot from the “direct side” or from the “equicorrelation side,” there is no need for simulation (as earlier from the direct side, when implementing the direct estimator via Algorithm 1) or numerical optimization (as earlier from the equicorrelation side, when implementing the matching estimator via solution of the first-order conditions (22)-(23)). Instead, the trivially simple closed-form estimator (29)-(30) is always applicable.
@@
We earlier observed that the empirical $DMSE^{avg}_{R,N\infty}(k)$ signature plots for growth and inflation appeared identical. The corollary shows this is not a data-specific coincidence. Theoretically, the two curves must coincide for every $k$.
@@ 3.3 On the Assessment of Equicorrelation Model Fit
As we have seen, equicorrelation ultimately emerges as a device for convenient calculation and understanding of direct signature plots, rather than as a separate “model” producing separate signature plots. In particular, nothing we have done requires that the equicorrelation be “true.” Nevertheless, it may be of interest in some contexts to assess whether forecast errors are truly equicorrelated, which can be done (under stronger assumptions than those invoked thus far) by maximum-likelihood estimation of a dynamic-factor model, followed by likelihood-ratio tests of the equicorrelation restrictions.
@@
The DFM (31) is already in state-space form, and one pass of the Kalman filter yields the innovations needed for exact Gaussian likelihood evaluation and optimization, and it also accounts for missing observations associated with survey entry and exit. One may also impose the weak or strong equicorrelation restrictions (33) or (34), respectively, and assess them using likelihood-ratio tests. Unsurprisingly, such formal tests (not reported) reject equicorrelation – a stylized “toy” data-generating process if ever there was one – for both growth and inflation, so we will not pursue formal testing.
@@ 4 Summary, Conclusions, and Directions for Future Research
We have studied the properties of macroeconomic survey forecast response averages as the number of survey respondents grows, characterizing the speed and pattern of the “gains from diversification” and their eventual decrease with “portfolio size” (the number of survey respondents) in both (1) the key real-world data-based environment of the U.S. Survey of Professional Forecasters (SPF), and (2) the theoretical model-based environment of equicorrelated forecast errors. We proceeded by proposing and comparing various direct and model-based “crowd size signature plots,” which summarize the forecasting performance of $k$-average forecasts as a function of $k$, where $k$ is the number of forecasts in the average. We then estimated the equicorrelation model for growth and inflation forecast errors by choosing model parameters to minimize the divergence between direct and model-based signature plots.